# What makes a good feedforward computational graph?

### v1

@@ 1 Introduction
Figure 1: In this paper, we propose two measures which can be used to assess suitability of feedforward computational graphs for neural networks – mixing time (lower is better; Sec. 4) and (normalised) minimax fidelity (higher is better; Sec. 5). Here we plot these metrics as attained by various graph generators—at each point in a sequence, the number of nodes doubles, starting from $16$. Using FunSearch (Romera-Paredes et al., 2023), we discover the FS graphs, which have favourable $O(\operatorname{polylog}n)$ mixing time, while retaining higher fidelity compared to a fully connected graph (Sec. 6). Several graphs may be viewed in App. C.
@@ 2 Motivation
We were inspired to seek an answer to this question, given that there is already a rich, extensive body of work on studying undirected computational graphs (commonly referred to as graph rewiring). Therein, important issues such as oversmoothing (Li et al., 2018; Oono & Suzuki, 2019; Keriven, 2022), oversquashing (Alon & Yahav, 2021; Di Giovanni et al., 2024) and under-reaching (Barceló et al., 2020) have all been identified, and related to the input graph topology. In response, a substantial amount of methods have been proposed to modify the provided input graph for improved diffusion (Gasteiger et al., 2019), curvature (Topping et al., 2022; Fesser & Weber, 2024), effective resistance (Arnaiz-Rodríguez et al., 2022), or reducing smoothing (Azabou et al., 2023) and commute time (Sterner et al., 2024).
@@
Unfortunately, to the best of our knowledge, the state-of-the-art is not as rich in the domain of feedforward graphs. This includes the domain of mathematics, wherein many important concepts have not been generalised to the directed case as they may rely on spectral properties of the graph structure (Chung, 1997), and many spectral properties are ill-defined on a feedforward graph. In terms of practical usage, most of the heavy lifting is done either by fully connected feedforward graphs or locally-connected sliding-window graphs. And while recent work has identified limitations of high-indegree feedforward graphs through over-squashing (Barbero et al., 2024) and dispersion (Veličković et al., 2024), most such works do not actively offer a different computational graph structure, nor do they offer any principles that can be used to derive one. We seek to fill this gap.
@@ 3 Theoretical foundations
  • The locally connected feedforward graph, which draws an edge between all allowed pairs of nodes up to a distance $\kappa\geq 0$ apart: =$\mathcal{E}=\{(a,b)\ |\ b-\kappa\leq a\leq b\}$. A special case of $\kappa=1$ is known as a line graph. Indegrees and outdegrees are $\delta_{i\leftarrow}=\delta_{i\rightarrow}=\kappa+1$.
@@ 4.2 Stationary distributions
Certainly $\mathbf{1}_{\tau}$ is stationary. To see that it is unique, consider a distribution $\boldsymbol{\pi}$ other than $\mathbf{1}_{\tau}$. Let $i<n$ be its smallest vertex with $\pi_{i}\not=0$. When applying the random walk matrix $\mathbf{W}$, we can see that the probability of being at $i$ after one step is $\pi_{i}/\delta_{i\rightarrow}$. Since $\delta_{i\rightarrow}>0$, we deduce that $\boldsymbol{\pi}$ is not stationary. ∎
@@ 4.3 Mixing times
We call this quantity the averaged mixing time (Díaz et al., 2024). We will now demonstrate its behaviour is as expected on two commonly used graph distributions.
@@ 4.3.1 Line graphs
$\frac{1}{n}\sum_{i=0}^{n-1}\sum_{j<i}\left({t\atop i}\right)2^{-t}=\frac{1}{n}\sum_{i=0}^{n-1}(n-i-1)\left({t\atop i}\right)2^{-t}.$
@@
$\sum_{i=0}^{n-1}(n-1)\left({t\atop i}\right)2^{-t}=(n-1)$
$\sum_{i=0}^{n-1}i\left({t\atop i}\right)2^{-t}=\sum_{i=0}^{t}i\left({t\atop i}\right)2^{-t}=t/2$
@@
$\frac{1}{n}\left(n-1-\frac{t}{2}\right)=1-\frac{t+1}{2n}.$
This is less than $1/4$ when $t\geq 3n/2$. Hence, the averaged mixing time is $23n/$ 2, which is $O(n)$.
@@ 4.4 Discussion about mixing time as a good measure
Hence $\kappa\mathbf{W}$ is the adjacency matrix $\mathbf{A}$ for the graph. Mixing time measures how close $\mathbf{W}^{t}\mathbf{e}_{i}$ is to the stationary distribution. In other words, it measures the size of $\mathbf{W}^{t}_{n-1,i}$. This is $1/\kappa^{t}$ times the $a_{n-1,i}$ entry of $\mathbf{A}^{t}$. Now this entry of $\mathbf{A}^{t}$ counts the number of oriented paths of length $t$ in $\mathcal{G}$ from vertex $i$ to the terminal vertex. So, we have the conjectured equivalence: mixing time is small (e.g. $O(\log n)$) if and only if there are ‘lots’ of paths from ‘most’ vertices to $\tau$. As having many paths seems useful for efficient data propagation, mixing time seems a good measure of how efficient our network will be under this computational graph.
@@ 5.2.2 Line graphs
This quantity also decays to zero as $n\rightarrow\infty$ (see Appendix A for a proof using Stirling’s approximation), but its normalised minimax fidelity consistently grows as $\sqrt{n/\pi}$ (see Fig. 1—indeed, line graphs have the highest initial fidelity of all considered graphs).
@@ 6 The FunSearch (FS) graph generator
There are two core issues with this graph left to fix from an asymptotics point of view. First, the indegree of each node is $O(n)$ as $n\rightarrow\infty$, which is undesirable. Second, the graph does not exhibit self-similarity: it will only work well if the predictions will take place in the final node block, which will be exposed to the “final triangle” of full edges.
@@ 6.1 Analysing the mixing time of the FS graph
For reasons of brevity, we provide a full proof of this Theorem (along with supplementary remarks concerning how to construct the graph to match the assumptions of the Theorem) in Appendix B.
@@ 7 Results
For details on the models’ architectures and optimisation, please see Appendix D. As one can see in Figure 3, the FS graph leads in general to the best performance in distribution and best out of distribution generalisation when compared to other sparse graphs—matching the intuition found by our metrics. More strongly, in the Parity task the FS graph manages to effectively match the fully connected graph, while having a substantially lower number of edges, and while its performance tends to be similar to the Poisson($0.8$) graph inthe Maximum task, in the Second Maximum and Parity tasks there’s a significant gap in performance between the FunSearch graph and all the other sparse graphs.
@@ 8 Conclusions
In this work, we have embarked on a detailed study of feedforward computational graphs, attempting to chart a novel path that could enable for a principled discovery of useful feedforward blueprints. While it is apparent that analysing these graphs is substantially less straightforward than doing so with their undirected counterparts, we believe the outcomes to have been fruitful. Namely, we proposed two well-justified metrics of feedforward information propagation, used them to automatically discover an interesting graph generator, and demonstrated its strong performance on several carefully crafted benchmarks. We are hopeful that our work will inspire targeted follow-up studies in this space, which may usher in a new paradigm of graph rewiring.
@@ Acknowledgements
We would like to extend deep thanks to Steven Kapturowski for detailed assistance with the parity experiment, as well as Alexander Novikov for kindly supporting us as we were setting up our FunSearch experiments. We are also grateful to Razvan Pascanu and Simon Osindero for diligently reviewing this paper and providing valuable comments.
@@
This paper presents a novel framework for improving the information propagation quality in feedforward models. Such models are highly prevalent, and innovations in this space could eventually lead to improvements in model sparsity (allowing for serving more powerful models on edge devices, for example) and model out-of-distribution generalisation (alowing more capable reasoning and scientific advances). Any societal consequences of our work could hence be equated to the societal consequences of accelerating these flavours of research.
Appendix A Proof that minimax fidelity in a line graph decays to zero with increasing size
@@
$\lim_{k\to\infty}\frac{\binom{2k}{k}}{2^{2k}}=\lim_{k\to\infty}\frac{2k!}{k!^{2}}\cdot\frac{1}{2^{2k}}=$
@@
Appendix B Proof of Theorem 6.1.
@@
Since we are mainly interested in the asymptotic number of depth levels, we make a simplifying assumption that both $\log n$ is an integer and that $n$ is divisible by each power of $\log n$. In this situation, the total number of levels $D$ is such that $n=\log^{D}n$. This implies $D=\log_{\log n}n=\frac{\log n}{\log\log n}$, i.e., the number of depth levels is $O\left(\frac{\log n}{\log\log n}\right)$.
With this in mind, we aim to quantify $T\left(\frac{\log n}{\log\log n},\log n\right)$, once again under the simplifying assumption that all of these variables are integers. To do this, we will establish several upper bounds on $T(d,k)$, assuming pessimistic behaviour from the random walker.
@@
Then, once the walker hits the final block in its current depth level ($k=0$), it automatically transitions into the next level, which is further subdivided into $\log n$ blocks. In the worst-case scenario, the walker will have landed in the very first block of the next depth level, and will need to traverse them all. Hence, $T(d,0)\leq T(d-1,\log n)$ for all $d>0$.
@@
We can now solve for all of these upper bounds as if they are equalities, yielding the following recurrence relations:
$T(d,k)$ $=1+\alpha T(d,k-1)+(1-\alpha)T(d,k)$ (7)
$T(d,0)$ $=T(d-1,\log n)$ (8)
$T(0,k)$ $=0$ (9)
The key observation is, since $\alpha$ is assumed identical in every depth level, and $k$ is always reset to $\log n$ after each depth level, we can equivalently represent the total walk time as a direct sum of walk times on each depth level.
Then, since each depth level is represented as a chain of $\log n$ states, $T(d,k)$ may be observed as the expected time of traversing a line graph of $k+(d-1)\log n$ steps and transition probability $\alpha$. The total number of such steps is $D\log n$, for $D$ being the number of depth levels.
@@
$\sum_{j\leq D\log n}\left({t\atop j}\right)\alpha^{j}(1-\alpha)^{t-j}<1/4.$
For ease of notation, set $N=D\log n$. The left hand side is the cumulative distribution function for the binomial random variable $X(t,\alpha)$. So we must find the smallest value of $t$ such that $\mathbb{P}(X(t,\alpha)\leq N)<1/4$. Now, Hoeffding’s inequality gives that
@@
When $\alpha$ is fixed and $N$ is large, this is certainly less than $1/4$. This establishes that the mixing time is at most $O(N)=O(D\log n)$. Note that when $t=D\log n$, then
$\sum_{j\leq D\log n}\left({t\atop j}\right)\alpha^{j}(1-\alpha)^{t-j}=1.$
So the mixing time is certainly at least $D\log n$.
Therefore, plugging in the appropriate values of $D$, we can conclude that the final mixing time will be $O\left(\frac{\log^{2}n}{\log\log n}\right)$, which is $O(\operatorname{polylog}n)$, completing the proof. ∎
Remark B.1.
@@
Since there will be $\frac{\log n}{\log\log n}$ depth levels in total, we cannot maintain a fixed outdegree of $\delta_{i\rightarrow}$ at each depth level, as the ratio is then $\frac{\delta_{i\rightarrow}}{\delta_{i\rightarrow}\frac{\log n}{\log\log n}}=\frac{\log\log n}{\log n}$, which will decay to zero as $n\rightarrow\infty$ and cannot be lower-bounded by a constant. This implies that the relative out-degree of each node in each block must decay with increasing depth level.
@@
Appendix C Adjacency matrix gallery

### v2

@@ 1 Introduction
Figure 1: In this paper, we propose two measures which can be used to assess suitability of feedforward computational graphs for neural networks – mixing time (lower is better; Sec. 4) and (normalised) minimax fidelity (higher is better; Sec. 5). Here we plot these metrics as attained by various graph generators—at each point in a sequence, the number of nodes doubles, starting from $16$. Using FunSearch (Romera-Paredes et al., 2023), we discover the FS graphs, which have favourable $O(\operatorname{polylog}n)$ mixing time, while retaining higher fidelity compared to a fully connected graph (Sec. 6). Several of these graphs may be viewed in Appendix D.
@@
Note that by “feedforward” we imply a statement about our input data—not the neural network processing it! That is, the assumption that our input nodes have a sequential ordering to them which must be respected when processing them—see Section 3. We therefore do not design our framework with more general feedforward neural networks (such as MLPs) in mind—however, we believe some of the ideas explored here could be used to study propagation in such unrestricted networks as well.
@@ 2 Motivation
We were inspired to seek an answer to this question, given that there is already a rich, extensive body of work on studying undirected computational graphs (commonly referred to as graph rewiring). Therein, important issues such as oversmoothing (Li et al., 2018; Oono & Suzuki, 2019; Keriven, 2022), oversquashing (Alon & Yahav, 2021; Di Giovanni et al., 2024) and under-reaching (Barceló et al., 2020) have all been identified, and related to the input graph topology. In response, a substantial amount of methods have been proposed to modify the provided input graph for improved diffusion (Gasteiger et al., 2019), curvature (Topping et al., 2022; Fesser & Weber, 2024), effective resistance (Arnaiz-Rodríguez et al., 2022), or reducing smoothing (Azabou et al., 2023) and commute time (Sterner et al., 2024). While efforts were made to generalise these techniques to directed graphs (Maskey et al., 2023), as well as for sparsifying attention-based architectures (Zaheer et al., 2020; Liao & Póczos, 2024), in all previous cases they focussed on graphs where backwards edges are explicitly allowed.
@@
Unfortunately, to the best of our knowledge, the state-of-the-art is not as rich in the domain of feedforward graphs. This includes the domain of mathematics, wherein many important concepts have not been generalised to the directed case [^1] We are aware of one prior work generalising undirected expanders to the concept of a feedforward extender (Csóka & Grabowski, 2022)—we leave making any connections of extender graphs to the propagation properties studied here for future work. as they may rely on spectral properties of the graph structure (Chung, 1997), and many spectral properties are ill-defined on a feedforward graph. In terms of practical usage, most of the heavy lifting is done either by fully connected feedforward graphs or locally-connected sliding-window graphs. And while recent work has identified limitations of high-indegree feedforward graphs through over-squashing (Barbero et al., 2024) and dispersion (Veličković et al., 2024), most such works do not actively offer a different computational graph structure, nor do they offer any principles that can be used to derive one. We seek to fill this gap.
@@ 3 Theoretical Foundations
  • The locally connected feedforward graph, which draws an edge between all allowed pairs of nodes up to a distance $\kappa\geq 0$ apart: =$\mathcal{E}=\{(a,b)\ |\ b-\kappa\leq a\leq b\}$. A special case of $\kappa=1$ is known as a line graph. Indegrees and outdegrees are $\delta_{i\leftarrow}=\delta_{i\rightarrow}=\kappa+1$ provided $\kappa\leq i\leq n-\kappa-1$.
@@ 4.2 Stationary Distributions
Certainly $\mathbf{1}_{\tau}$ is stationary. To see that it is unique, consider a distribution $\boldsymbol{\pi}$ other than $\mathbf{1}_{\tau}$. Let $i<n-1$ be its smallest vertex with $\pi_{i}\not=0$. When applying the random walk matrix $\mathbf{W}$, we can see that the probability of being at $i$ after one step is $\pi_{i}/\delta_{i\rightarrow}$. Since $\delta_{i\rightarrow}>1$, we deduce that $\boldsymbol{\pi}$ is not stationary. ∎
@@ 4.3 Mixing Times
We call this quantity the averaged mixing time (Espuny Díaz et al., 2024). We will now demonstrate its behaviour is as expected on two commonly used graph distributions.
@@ 4.3.1 Line Graphs
$\frac{1}{n}\sum_{i=0}^{n-1}\sum_{j<i}\binom{t}{j}2^{-t}=\frac{1}{n}\sum_{i=0}^{n-1}(n-i-1)\binom{t}{i}2^{-t}.$
@@
$\sum_{i=0}^{n-1}(n-1)\binom{t}{i}2^{-t}=(n-1)$
$\sum_{i=0}^{n-1}i\binom{t}{i}2^{-t}=\sum_{i=0}^{t}i\binom{t}{i}2^{-t}=t/2$
@@
$\frac{1}{n}\left(n-1-\frac{t}{2}\right)=1-\frac{t+2}{2n}.$
This is less than $1/4$ when $t>3n/2$. Hence, the averaged mixing time is $23n/$ 2, which is $O(n)$.
@@ 4.4 Discussion About Mixing Time as a Good Measure
Hence $\kappa\mathbf{W}$ is the adjacency matrix $\mathbf{A}$ for the graph. Mixing time measures how close $\mathbf{W}^{t}\mathbf{e}_{i}$ is to the stationary distribution. In other words, it measures the size of $\mathbf{W}^{t}_{n-1,i}$. This is $1/\kappa^{t}$ times the $a_{n-1,i}$ entry of $\mathbf{A}^{t}$. Now this entry of $\mathbf{A}^{t}$ counts the number of oriented paths of length $t$ in $\mathcal{G}$ from vertex $i$ to the terminal vertex. So, we have the conjectured equivalence: mixing time is small (e.g. $O(\log n)$) if and only if there are ‘lots’ of paths from ‘most’ vertices to $\tau$. In fact, we can formally prove the following proposition:
Proposition 4.2.
Suppose that the outdegree of every vertex other than $\tau$ is at least $2$. Let $t$ be the average mixing time. Then for some $s\leq t$, the average number of paths from vertex $i$ to $\tau$ with length $s$ is at least $(3/4t)2^{s}$, where the average is taken over all vertices $i$ between 00 and $n-1$.
One way to interpret this is for the mixing time of $\mathcal{G}$ to be ‘small’, then there must be exponentially many paths from a typical vertex to $\tau$ with length less than the mixing time. We prove this statement in Appendix A. As having many paths seems useful for efficient data propagation, mixing time seems a good measure of how efficient our network will be under this computational graph.
@@ 5.2.2 Line Graphs
This quantity also decays to zero as $n\rightarrow\infty$ (see Appendix B for a proof using Stirling’s approximation), but its normalised minimax fidelity consistently grows as $\sqrt{n/\pi}$ (see Figure 1—indeed, line graphs have the highest initial fidelity of all considered graphs).
@@ 6 The FunSearch (FS) Graph Generator
There are two core issues with this graph left to fix from an asymptotics point of view. First, the indegree of each node is $O(n/\log n)$ as $n\rightarrow\infty$, which is undesirable. Second, the graph does not exhibit self-similarity: it will only work well if the predictions will take place in the final node block, which will be exposed to the “final triangle” of full edges.
@@ 6.1 Analysing the Mixing Time of the FS Graph
For reasons of brevity, we provide a full proof of this Theorem (along with supplementary remarks concerning how to construct the graph to match the assumptions of the Theorem) in Appendix C.
@@ 7 Results
For details on the models’ architectures and optimisation, please see Appendix E. As one can see in Figure 3, the FS graph leads in general to the best performance in distribution and best out of distribution generalisation when compared to other sparse graphs—matching the intuition found by our metrics. More strongly, in the Parity task the FS graph manages to effectively match the fully connected graph, while having a substantially lower number of edges, and while its performance tends to be similar to the Poisson($0.8$) graph inthe Maximum task, in the Second Maximum and Parity tasks there’s a significant gap in performance between the FunSearch graph and all the other sparse graphs.
@@
On the Importance of Self-Edges and In-Degrees
Our theoretical framework always assumes that all self-edges are included, as a valuable way to preserve each node’s information over time. Accordingly, removing it led to significant performance regressions on all tasks – especially on the Parity task where multiple elements need to interact meaningfully for the final answer computation.
We also evaluated our graph generators at different orders-of-magnitude of in-degree – from $O(1)$ to $O(\log n)$ to $O(\sqrt{n})$. We find that, as the in-degrees are increased, the models get more performant in-distribution, without a significant effect on overall out-of-distribution performance.
@@ 8 Conclusions
In this work, we have embarked on a detailed study of feedforward computational graphs, attempting to chart a novel path that could enable for a principled discovery of useful feedforward blueprints. While it is apparent that analysing these graphs is substantially less straightforward than doing so with their undirected counterparts, we believe the outcomes to have been fruitful. Namely, we proposed two well-justified metrics of feedforward information propagation, used them to automatically discover an interesting graph generator, and demonstrated its strong performance on several carefully crafted benchmarks.
We are hopeful that our work will inspire targeted follow-up studies in this space, which may usher in a new paradigm of graph rewiring. To this end, we particularly believe it would be worthwhile to study computational graphs that dynamically change across layers – as opposed to a single fixed graph studied here – as well as attempting to establish a more formal framework connecting mixing time, fidelity and similar metrics. The framework of Lim et al. (2024) could be useful for reasoning about feedforward computational graphs spanning multiple layers.
@@ Acknowledgements
We would like to extend deep thanks to Steven Kapturowski for detailed assistance with the parity experiment, as well as Alexander Novikov for kindly supporting us as we were setting up our FunSearch experiments. We are also grateful to Razvan Pascanu and Simon Osindero for diligently reviewing this paper and providing valuable comments. Marc Lackenby was partially supported by the Engineering and Physical Sciences Research Council (grant number EP/Y004256/1).
@@
This paper presents a novel framework for improving the information propagation quality across data with a feedforward constraint. Such models are highly prevalent, and innovations in this space could eventually lead to improvements in model sparsity (allowing for serving more powerful models on edge devices, for example) and model out-of-distribution generalisation (alowing more capable reasoning and scientific advances). Any societal consequences of our work could hence be equated to the societal consequences of accelerating these flavours of research.
Appendix A Proof of Proposition 4.2.
Proof.
For $10\leq i<n-$ 1, let $p(i,s)$ be the number of paths from vertex $i$ to $\tau$ with length $s$. Then
$(\mathbf{W}^{t})_{\tau i}\leq\sum_{s\leq t}\frac{p(i,s)}{2^{s}}.$
This is because $(\mathbf{W}^{t})_{\tau i}$ represents the probability of a random walk starting at $i$ and reaching $\tau$ by time $t$. This is equal to the sum, over all $s\leq t$, of the probability that the random walk first reaches $\tau$ at time $s$. For any such $s$, this probability is equal to the sum, over all paths from $i$ to $\tau$ with length $s$ avoiding the loop based at $\tau$, of the probability of taking that path. This latter probability is at most $1/2^{s}$ because at each step along the path there were at least two options that could have been taken.
For $t$ equal to the averaged mixing time,
$\frac{1}{n}\sum_{i=0}^{n-1}(\mathbf{W}^{t})_{\tau i}\geq\frac{3}{4}.$
So,
$\sum_{s\leq t}\frac{1}{n}\sum_{i=0}^{n-1}\frac{p(i,s)}{2^{s}}=\frac{1}{n}\sum_{i=0}^{n-1}\sum_{s\leq t}\frac{p(i,s)}{2^{s}}\geq\frac{1}{n}\sum_{i=0}^{n-1}(\mathbf{W}^{t})_{\tau i}\geq\frac{3}{4}.$
Therefore, for some $s\leq t$,
$\frac{1}{n}\sum_{i=0}^{n-1}\frac{p(i,s)}{2^{s}}\geq\frac{3}{4t}.$
So, the average number of paths from vertex $i$ to $\tau$ with length $s$ is at least $(3/4t)2^{s}$. ∎
Appendix B Proof that Minimax Fidelity in a Line Graph Decays to Zero with Increasing Size
@@
$\lim_{k\to\infty}\frac{\binom{2k}{k}}{2^{2k}}=\lim_{k\to\infty}\frac{(2k)!}{k!^{2}}\cdot\frac{1}{2^{2k}}=$
@@
Appendix C Proof of Theorem 6.1.
@@
Denoting the total number of levels by $D$, we have that
$n/(\lceil\log n\rceil)^{D+1}<1,\qquad n/(\lceil\log n\rceil)^{D}\geq 1$
and hence
$\log n<(D+1)\log\lceil\log n\rceil,\qquad\log n\geq D\log\lceil\log n\rceil.$
So
$D=\left\lfloor\frac{\log n}{\log\lceil\log n\rceil}\right\rfloor.$
With this in mind, we aim to quantify $T\left(\left\lfloor\frac{\log n}{\log\lceil\log n\rceil}\right\rfloor,\lceil\log n\rceil\right)$. To do this, we will establish several upper bounds on $T(d,k)$, assuming pessimistic behaviour from the random walker.
@@
Then, once the walker hits the final block in its current depth level ($k=0$), it automatically transitions into the next level, which is further subdivided into $\lceil\log n\rceil$ blocks. In the worst-case scenario, the walker will have landed in the very first block of the next depth level, and will need to traverse them all. Hence, $T(d,0)\leq T(d-1,\lceil\log n\rceil)$ for all $d>0$.
@@
We now define $\tilde{T}(d,k)$ recursively using the following recurrence relations:
$\tilde{T}(d,k)$ $=1+\alpha\tilde{T}(d,k-1)+(1-\alpha)\tilde{T}(d,k)$ (7)
$\tilde{T}(d,0)$ $=\tilde{T}(d-1,\lceil\log n\rceil)$ (8)
$\tilde{T}(0,k)$ $=0$ (9)
A simple induction gives that $T(d,k)\leq\tilde{T}(d,k)$. So it suffices to find an upper bound for $\tilde{T}(d,k)$.
The key observation is, since $\alpha$ is assumed identical in every depth level, and $k$ is always reset to $\lceil\log n\rceil$ after each depth level, we can equivalently represent the total walk time as a direct sum of walk times on each depth level.
Then, since each depth level is represented as a chain of $\lceil\log n\rceil$ states, $\tilde{T}(d,k)$ may be observed as the expected time of traversing a line graph of $k+(d-1)\lceil\log n\rceil$ steps and transition probability $\alpha$. The total number of such steps is $D\lceil\log n\rceil$, for $D$ being the number of depth levels.
@@
$\sum_{j\leq D\lceil\log n\rceil}\binom{t}{j}\alpha^{j}(1-\alpha)^{t-j}<1/4.$
For ease of notation, set $N=D\lceil\log n\rceil$. The left hand side is the cumulative distribution function for the binomial random variable $X(t,\alpha)$. So we must find the smallest value of $t$ such that $\mathbb{P}(X(t,\alpha)\leq N)<1/4$. Now, Hoeffding’s inequality gives that
@@
When $\alpha$ is fixed and $N$ is large, this is certainly less than $1/4$. This establishes that the mixing time is at most $O(N)=O(D\log n)$. Note that when $t=D\lceil\log n\rceil$, then
$\sum_{j\leq D\lceil\log n\rceil}\binom{t}{j}\alpha^{j}(1-\alpha)^{t-j}=1.$
So the mixing time is certainly at least $D\lceil\log n\rceil$.
Therefore, plugging in the appropriate values of $D$, we can conclude that the final mixing time will be $O\left(\frac{\log^{2}n}{\log\log n}\right)$, which is $O(\mathrm{polylog}\,n)$, completing the proof. ∎
Remark C.1.
@@
Since there will be approximately $\frac{\log n}{\log\log n}$ depth levels in total, we cannot maintain a fixed outdegree of $\delta_{i\rightarrow}$ at each depth level, as the ratio is then $\frac{\delta_{i\rightarrow}}{\delta_{i\rightarrow}\frac{\log n}{\log\log n}}=\frac{\log\log n}{\log n}$, which will decay to zero as $n\rightarrow\infty$ and cannot be lower-bounded by a constant. This implies that the relative out-degree of each node in each block must decay with increasing depth level.
@@
Appendix D Adjacency Matrix Gallery